# Lost in Tracking Translation: A Comprehensive Analysis of Visual SLAM in Human-Centered XR and IoT Ecosystems

### v1

@@ 3.2.1. Environmental Challenges
Finally, the ability of the sensors to accurately sense the environmental elements adds another layer of complexity. Sensors’ abilities to sense the environment vary based on their specifications, such as the frame rate they can support, their focal length, and sensor sizes. The sensor’s capabilities can vary across application scenarios and impact the tracking accuracy. In addition, even if two sensors are the same based on all the metrics mentioned above, the sensor can have variations due to manufacturing defects or natural variations in the materials used to synthesize them (Chong et al., 2015; Servières et al., 2021; Andreopoulos and Tsotsos, 2011).
@@ 3.2.2. Locomotion-based Challenges
Another key challenge is scaling tracking algorithms to expanded maps and complex environments, which requires significant computational resources. For example, Loop Closure Detection requires identifying previously visited locales, which scales as the environment expands (Latif et al., 2013). These issues are exacerbated in XR environments that pose the additional challenges of constrained resources (Semenova et al., 2022). Additionally, XR headsets often have multiple cameras, which introduces synchronization and calibration challenges (Yu et al., 2021b; Saeedi et al., 2016; Halmetschlager-Funek et al., 2018). These issues impact the quality and consistency of the available data to the algorithms and the data used to train the models. The performance significantly degrades if input data distribution significantly shifts. Also, the robustness of these methods may vary across application and environment scenarios. Finally, unlike traditional SLAM methods, learning-based methods are black boxes and offer little to no interpretability. For reliable tracking, it is crucial to understand and solve multifaceted challenges for applications, including autonomous driving, robotics, and mixed reality.
@@ 4.1.1. State-of-the-Art Tracking Algorithms
In our analysis, DeepVO (Wang et al., 2017) is the only pure deep learning-based end-to-end VO system, and it does not rely even on traditional keyframe detection or explicit feature extraction.
@@ 4.1.2. Datasets and Application Scenarios
We select three datasets that map to the three sample combinations of the environment, object, and motion, shown in Figure 1. While there can be other possible combinations, these three use cases represent the most common tracking applications and the most widely used datasets for these applications.
@@ 4.1.3. Performance Metrics for Evaluation
  • Relative Pose Error (RPE) measures the difference between relative transformations at time instances $i$ and $i+k$ for different values of $k$. If $\Delta T_{gt}(i,i+k)$ and $\Delta T_{est}(i,i+k)$ are the relative transformations for ground truth and estimated trajectories respectively, then:
@@ 4.2. Head-to-Head Performance Comparison at All Levels
2 – Dataset-Level Question: Do tracking methods show robustness to variations in dataset characteristics? To explore this, we analyzed the performance of algorithms across three distinct datasets, as shown in Table 6. Interestingly, the best-performing algorithm was not the same across the datasets. For the KITTI dataset, ORB-SLAM3 (S) emerged as the strongest, performing best on 5 of the 11 sequences, while KP3D followed closely with 3 sequences. However, when applied to the HoloSet dataset, ORB-SLAM3 (S) failed completely, unable to generate trajectories for any sequence. This pattern was reinforced by examining ATE values and the coefficient of variation (CoV), indicating that strong performance on one dataset does not necessarily translate to others, particularly when faced with different environmental conditions, motion dynamics, and object types.
We also visually demonstrate the performance of a sample algorithm across sequences from different datasets. In Figure 3, TartanVO’s performance is illustrated across three datasets, showing significant variation across sequences within each dataset. While it performs reasonably well in the EuRoC MH02 sequence with structured, textured surfaces, its accuracy drops notably in KITTI’s outdoor high-speed sequence, where low texture and fast motion introduce more complexity. The HoloSet suburban walk sequence further highlights this inconsistency, where TartanVO struggles with scale estimation and trajectory alignment due to the absence of loop closure detection.
@@
3 – Sequence-Level Question: Do algorithms show consistent performance across sequences within a single dataset? We analyzed the results across sequences to explore this, focusing on the data in Table 3 and Table 6. Among the tracking methods, DeepVO demonstrates the most consistent performance, with the lowest coefficient of variation (CoV) across sequences at 0.13. Its average ATE of 1.77m is relatively low, positioning it among the better-performing methods despite not being the top performer for any sequence. In contrast, top-performing methods like KP3D and ORB-SLAM3 (S) exhibit higher CoV values of 0.58 and 0.85, indicating more variability in performance across sequences.
@@
4 – Sample-level Question: Can the methods handle the variations in the samples within sequence without sample-level fine-tuning? Despite the good performance of some visual SLAM methods under “ideal conditions”, there is no clear winner on method, dataset, and sequence levels. Our evaluation has unveiled several algorithmic, environmental, and locomotion challenges that SLAM methods face on different levels of the pipeline. To investigate further, we evaluate sample-level attributes. We took six samples from EuRoC sequences (MH01, MH03, V203): one baseline vanilla sample and five samples with common visual features essential for data characterization in SLAM (Ali and Zhang, 2022; Saeedi et al., 2019). These features include occlusion, lighting, reflective surface, image motion blur, and texture variability. We tracked the feature of each sample with VINS-Fusion (Qin and Shen, 2018) feature extractor as shown in Figure 4.
We observe that the number of features ranges from 127 per frame to a mere 17 features per frame. While certain attributes might favor the algorithms, leading to reliable tracking (as shown in Figure 4(a)), others can significantly strain the system, as seen in Figure  4(f). Any sudden movement (causing blurry image), low lighting, low textured environment, or occlusions can hinder the methods’ view of certain features. Whereas shiny reflective surfaces can create false features, compromising tracking accuracy. A method’s ability to consistently recognize and adapt to these varying attributes without manual intervention or fine-tuning signifies its robustness.
Key Takeaway. The disparity in results across the sequences, samples, and visual features indicates that achieving a truly generalized method for handling the broad spectrum of sample attributes remains an open question. A robust and domain-adaptive SLAM can be developed by understanding data at various levels, from method down to sample level.
@@ 4.3. Impact of Choices in Visual SLAM Components and Pipelines
Deep Learning Methods. The deep learning-based method, DeepVO, stands out for its consistency across sequences within a single dataset, demonstrated by the lowest CoV of 0.13 on EuRoC. However, its end-to-end performance in terms of average ATE is not as competitive, with values such as 1.77m on EuRoC and 37.94m on HoloSet, indicating that while it performs consistently, its overall error is higher than most hybrid and traditional approaches. Additionally, deep learning methods tend to struggle on datasets with more complex environmental or motion variations, as indicated by the large ATE on HoloSet.
@@ 4.4. Performance in Tackling Environmental, Locomotion, and Algorithmic Challenges
To compare how SLAM, deep learning (DL), and hybrid methods tackle environmental, locomotion, and algorithmic challenges, we analyze their performance across the datasets (EuRoC, KITTI, and HoloSet) summarized in Table 6, in addition to the individual results for datasets. These datasets capture a range of conditions, from indoor to outdoor environments, controlled settings to dynamic scenes, and varying frame rates and resolutions. Each method’s ability to adapt to these challenges reveals insights into the strengths and weaknesses of different computational approaches.
Traditional Methods. SLAM methods like ORB-SLAM3 (S) and ORB-SLAM3 (M) perform well in controlled environments but struggle as environmental and motion complexity increases. On EuRoC’s easier sequences, such as MH01, ORB-SLAM3 (S) achieves a low ATE of 0.24m, aided by EuRoC’s compact resolution and controlled environment, which facilitates smoother loop closures and reduces scalability challenges. However, as sequences become more turbulent (e.g., V202), ORB-SLAM3 (S) ’s ATE increases to 0.84m, reflecting the limitations of traditional methods in more complex indoor environments. In KITTI, while SLAM methods benefit from high-quality calibration and ground truth data, broader variability in lighting and environmental factors challenges their robustness. For instance, ORB-SLAM3 (S) records an ATE of 4.80m in sequence 00, but as environmental complexity increases in sequence 02, the ATE jumps to 11.2m. This highlights SLAM’s struggle with outdoor settings where rapid changes in lighting and environment occur. On HoloSet, SLAM methods fail entirely, and they are unable to handle the unpredictable, multi-directional motion of the human-controlled XR headset. Both ORB-SLAM3 (S) and ORB-SLAM3 (M) are unable to produce valid trajectories, reflecting their difficulty in handling high frame rates and the uncontrolled dynamics of human motion. The lack of fine-tuning for HoloSet’s novel environmental conditions and higher frame rates further exacerbates their poor performance.
@@
Deep Learning Methods. DL methods exhibit consistency across sequences but generally incur higher error rates, especially in environments with significant variability. While DL methods provide consistent performance, their ATE values are higher than hybrid methods. For instance, in MH01, DeepVO records an ATE of 1.67m – consistent but less accurate compared to KP3D ’s 0.21m. This indicates that DL methods are less precise in handling turbulent motion within controlled indoor environments. On KITTI, DL methods struggle more with outdoor complexity. For example, DeepVO cannot produce results for most sequences (denoted by “–” in the table), highlighting the difficulty of applying DL methods to rapidly changing outdoor environments, particularly when lighting conditions and scene complexity fluctuate. On the HoloSet dataset, DL methods perform better than SLAM but still lag behind hybrid methods. For example, DeepVO records an ATE of 33.0m in suburb-jog-seq1, significantly higher than hybrid methods like KP3D. While DL methods can handle unpredictable human motion, their error rates remain higher than hybrid methods, reflecting their limitations in highly dynamic, human-centered environments.
@@ 4.5. Effect of Human Factors in XR
Deep Learning Methods. DL methods, such as DeepVO, also show an ability to track human motion better than SLAM methods but with less precision than hybrid approaches. In HoloSet’s suburb-jog-seq1, DeepVO records an ATE of 33.0m, significantly higher than KP3D ’s performance. While DL methods benefit from their inherent flexibility in handling varying motions, their higher error rates suggest they are less adept at managing rapid, unpredictable changes in orientation and environment caused by human interaction in XR scenarios.

### v2

@@ 3.2.1. Environmental Challenges
Finally, the ability of the sensors to accurately sense the environmental elements adds another layer of complexity. Sensors’ abilities to sense the environment vary based on their specifications, such as the frame rate they can support, their focal length, and sensor sizes. The sensor’s capabilities can vary across application scenarios and impact the tracking accuracy. In addition, even if two sensors are the same based on all the metrics mentioned above, the sensor can have variations due to manufacturing defects or natural variations in the materials used to synthesize them (Chong et al., 2015; Servières et al., 2021; Andreopoulos and Tsotsos, 2011). Sensor limitations are an important component of environmental challenges. Hardware properties such as shutter type (rolling vs. global), resolution, field-of-view (FoV), frame rate, dynamic range, and sensor noise fundamentally constrain what tracking methods can achieve. For example, rolling shutter distortion during fast motion introduces geometric artifacts that degrade pose estimation unless explicitly modeled or corrected. Narrow FoV or low resolution reduces feature coverage and robustness, particularly in texture-poor scenes. A low frame rate exacerbates motion blur under fast locomotion, while a high frame rate increases computational demands. Finally, we note that sensor-related limitations—such as rolling shutter effects, limited field-of-view, and sensor noise, interact with environmental and locomotion factors to impact tracking accuracy; we analyze these effects in detail in Section 5.2.
@@ 3.2.2. Locomotion-based Challenges
Another key challenge is scaling tracking algorithms to expanded maps and complex environments, which requires significant computational resources. For example, Loop Closure Detection requires identifying previously visited locales, which scales as the environment expands (Latif et al., 2013). These issues are exacerbated in XR environments that pose the additional challenges of constrained resources (Semenova et al., 2022). Additionally, XR headsets often have multiple cameras, which introduces synchronization and calibration challenges (Yu et al., 2021b; Saeedi et al., 2016; Halmetschlager-Funek et al., 2018). These issues impact the quality and consistency of the available data to the algorithms and the data used to train the models. The performance significantly degrades if the input data distribution significantly shifts. Also, the robustness of these methods may vary across applications and environmental scenarios. Finally, unlike traditional SLAM methods, learning-based methods are black boxes and offer little to no interpretability. For reliable tracking, it is crucial to understand and solve multifaceted challenges for applications, including autonomous driving, robotics, and mixed reality.
@@ 4.1.1. State-of-the-Art Tracking Algorithms
In our analysis, DeepVO (Wang et al., 2017) is the only pure deep learning-based end-to-end VO system, and it does not rely on traditional keyframe detection or explicit feature extraction.
Evaluation Pipeline Note. To ensure consistency across all methods, we adopt a unified evaluation pipeline: all trajectories are processed through the same post-alignment and error computation steps (ATE and RPE). We use the publicly available open-source versions of each SLAM/VO method, with default configurations unless otherwise specified. Importantly, we do not apply method-specific tuning, multiple reruns, or per-sequence parameter optimization. As a result, our reported ATE/RPE may differ from those in the original papers, where such tuning or manual loop closure interventions were used to achieve best-case performance. This design choice ensures a fair and systematic comparison across methods, even if it exposes sensitivity to failure modes.
@@ 4.1.2. Datasets and Application Scenarios
We select three datasets that map to the three sample combinations of the environment, object, and motion, shown in Figure 1. While there can be other possible combinations, these three use cases represent the most common tracking applications and the most widely used datasets for these applications. We summarize the key sensor specifications for these datasets in Table 3.
Table 3. Sensor specifications for evaluated datasets.
Dataset Camera Type Resolution FPS Shutter Type FoV
KITTI Stereo grayscale 1392 $\times$ 512 10 Global $\sim$81
EuRoC Stereo monochrome 752 $\times$ 480 20 Global $\sim$60
HoloSet 4$\times$ grayscale + RGB + Depth Gray: 640 $\times$ 480 30 Rolling Gray: $\sim$86
RGB: 1920 $\times$ 1080 30 Rolling RGB: $\sim$81
Depth: 512 $\times$ 512 Variable (5-45) Rolling Depth: $\sim$75
@@
While there can be other possible combinations, these three use cases represent the most common tracking applications and the most widely used datasets for these applications. We acknowledge that other datasets such as ADVIO (Cortés et al., 2018), OxIOD (Chen et al., 2018), YTU (Gurturk et al., 2021), TUM VI (Schubert et al., 2018), ZJU-SenseTime (Jinyu et al., 2019), Oxford RobotCar (Maddern et al., 2017), UMA-VI (Zuñiga-Noël et al., 2020), PALVIO (Wang et al., 2022b), ICL-NUIM (Handa et al., 2014), and Málaga (Blanco-Claraco et al., 2014). Other includes datasets that focus on human gait (MAREA (Khandelwal and Wickström, 2017), OU-ISIR (Uddin et al., 2018)), occupancy (LARA (Niemann et al., 2020)), and activity recognition (USC-HAD (Zhang and Sawchuk, 2012), CMU-MMAC (De la Torre et al., 2009), Opportunity (Chavarriaga et al., 2013)) offer complementary challenges, including richer motion types, XR-relevant trajectories, eye tracking and sensor configurations such as visual-inertial streams. However, we restrict our study to datasets that provide visual-only streams to ensure consistency across the selected monocular, stereo, and learning-based SLAM/VO methods. Evaluating cross-modal SLAM methods, such as visual-inertial pipelines, is a promising direction for future work, but beyond the current paper’s scope.
@@ 4.1.3. Performance Metrics for Evaluation
  • Relative Pose Error (RPE) measures the difference between relative transformations at time instances $i$ and $i+k$ for different values of $k$. If $\Delta T_{gt}(i,i+k)$ and $\Delta T_{est}(i,i+k)$ are the relative transformations for ground truth and estimated trajectories, respectively, then:
@@ 4.2. Head-to-Head Performance Comparison at All Levels
2 – Dataset-Level Question: Do tracking methods show robustness to variations in dataset characteristics? To explore this, we analyzed the performance of algorithms across three distinct datasets, as shown in Table 7. Interestingly, the best-performing algorithm was not the same across the datasets. For the KITTI dataset, ORB-SLAM3 (S) emerged as the strongest, performing best on 5 of the 11 sequences, while KP3D followed closely with 3 sequences. However, when applied to the HoloSet dataset, ORB-SLAM3 (S) failed completely, unable to generate trajectories for any sequence. SLAM pipelines such as ORB-SLAM3 (S) rely on explicit loop closure and global optimization to maintain tracking consistency. However, human-controlled motion in XR environments (HoloSet) involves rapid, multi-directional head movements and limited scene revisitation, making loop closure ineffective and causing drift to accumulate rapidly. Moreover, HoloSet employs a rolling shutter camera, which introduces motion-induced artifacts during fast head rotations. These sensor characteristics, combined with fast angular motion, violate key assumptions of traditional SLAM systems, causing failure to initialize or maintain reliable tracking. SLAM methods tuned for global shutter and vehicle-like motion (EuRoC, KITTI) are fragile when exposed to a rolling shutter and rapid, 6-DoF head motion (HoloSet). This pattern was reinforced by examining ATE values and the coefficient of variation (CoV), indicating that strong performance on one dataset does not necessarily translate to others, particularly when faced with different environmental conditions, motion dynamics, and object types. In part, this is due to differences in sensor characteristics (see Section 5.2) and motion dynamics across datasets. For example, EuRoC and KITTI use global shutter cameras with relatively stable, smooth motion patterns (vehicle or drone-based). In contrast, HoloSet uses a rolling shutter camera and exhibits fast rotational motion from human head movements. SLAM methods tuned for global shutter and deep-learning methods trained on vehicle datasets often fail to generalize to these challenging XR motion conditions.
We also visually demonstrate the performance of a sample algorithm across sequences from different datasets. In Figure 3, TartanVO’s performance is illustrated across three datasets, showing significant variation across sequences within each dataset. While it performs reasonably well in the EuRoC MH02 sequence with structured, textured surfaces, its accuracy drops notably in KITTI’s outdoor high-speed sequence, where low texture and fast motion introduce more complexity. The HoloSet suburban walk sequence further highlights this inconsistency, where TartanVO struggles with scale estimation and trajectory alignment due to the absence of loop closure detection. Pure deep-learning methods like TartanVO perform inference based on temporal image sequences but do not explicitly model global consistency or geometric structure. As a result, they can generalize well to familiar motion patterns but exhibit instability when exposed to motion regimes not seen in training data. In particular, XR head motion includes abrupt changes in rotation and translation that differ significantly from the vehicle-based or drone-based motions typically present in SLAM training datasets. This mismatch leads to degraded performance, particularly in sequences with fast or unpredictable human movement.
@@
3 – Sequence-Level Question: Do algorithms show consistent performance across sequences within a single dataset? We analyzed the results across sequences to explore this, focusing on the data in Table 4 and Table 7. Among the tracking methods, DeepVO demonstrates the most consistent performance, with the lowest coefficient of variation (CoV) across sequences at 0.13. Its average ATE of 1.77m is relatively low, positioning it among the better-performing methods despite not being the top performer for any sequence. In contrast, top-performing methods like KP3D and ORB-SLAM3 (S) exhibit higher CoV values of 0.58 and 0.85, indicating more variability in performance across sequences. DL methods are less sensitive to scene texture but cannot handle rapid viewpoint changes or scale ambiguity. As a result, they produce consistent errors but often higher overall ATE, reflecting model bias toward learned motion patterns.
@@
4 – Sample-level Question: Can the methods handle the variations in the samples within a sequence without sample-level fine-tuning? Despite the good performance of some visual SLAM methods under “ideal conditions”, there is no clear winner on method, dataset, and sequence levels. Our evaluation has unveiled several algorithmic, environmental, and locomotion challenges that SLAM methods face on different levels of the pipeline. To investigate further, we evaluate sample-level attributes. We took six samples from EuRoC sequences (MH01, MH03, V203): one baseline vanilla sample and five samples with common visual features essential for data characterization in SLAM (Ali and Zhang, 2022; Saeedi et al., 2019). These features include occlusion, lighting, reflective surface, image motion blur, and texture variability. We tracked the feature of each sample with the VINS-Fusion (Qin and Shen, 2018) feature extractor as shown in Figure 4.
We observe that the number of features ranges from 127 per frame to a mere 17 features per frame. While certain attributes might favor the algorithms, leading to reliable tracking (as shown in Figure 4(a)), others can significantly strain the system, as seen in Figure 4(f). Any sudden movement (causing a blurry image), low lighting, low textured environment, or occlusions can hinder the methods’ view of certain features. Whereas shiny reflective surfaces can create false features, compromising tracking accuracy. A method’s ability to consistently recognize and adapt to these varying attributes without manual intervention or fine-tuning signifies its robustness.
Key Takeaway. The disparity in results across the sequences, samples, and visual features indicates that achieving a truly generalized method for handling the broad spectrum of sample attributes remains an open question. A robust and domain-adaptive SLAM can be developed by understanding data at various levels, from the method down to the sample level.
@@ 4.3. Impact of Choices in Visual SLAM Components and Pipelines
Deep Learning Methods. The deep learning-based method, DeepVO, stands out for its consistency across sequences within a single dataset, demonstrated by the lowest CoV of 0.13 on EuRoC. However, its end-to-end performance in terms of average ATE is not as competitive, with values such as 1.77m on EuRoC and 37.94m on HoloSet, indicating that while it performs consistently, its overall error is higher than most hybrid and traditional approaches. Additionally, deep learning methods tend to struggle on datasets with more complex environmental or motion variations, as indicated by the large ATE on HoloSet. The low CoV of DeepVO reflects its ability to produce stable predictions across sequences, but its relatively higher ATE indicates a systematic bias or model drift. Since DeepVO lacks explicit modeling of scale or 3D structure, its learned motion estimation is less sensitive to texture or loop closure but prone to accumulating global errors. As a result, while the method appears consistent, its absolute accuracy suffers, particularly in longer or more dynamic sequences.
@@ 4.4. Performance in Tackling Environmental, Locomotion, and Algorithmic Challenges
To compare how SLAM, deep learning (DL), and hybrid methods tackle environmental, locomotion, and algorithmic challenges, we analyze their performance across the datasets (EuRoC, KITTI, and HoloSet) summarized in Table 7, in addition to the individual results for the datasets. These datasets capture a range of conditions, from indoor to outdoor environments, controlled settings to dynamic scenes, and varying frame rates and resolutions. Each method’s ability to adapt to these challenges reveals insights into the strengths and weaknesses of different computational approaches.
Traditional Methods. SLAM methods like ORB-SLAM3 (S) and ORB-SLAM3 (M) perform well in controlled environments but struggle as environmental and motion complexity increases. On EuRoC’s easier sequences, such as MH01, ORB-SLAM3 (S) achieves a low ATE of 0.24m, aided by EuRoC’s compact resolution and controlled environment, which facilitates smoother loop closures and reduces scalability challenges. However, as sequences become more turbulent (e.g., V202), ORB-SLAM3 (S) ’s ATE increases to 0.84m, reflecting the limitations of traditional methods in more complex indoor environments. In KITTI, while SLAM methods benefit from high-quality calibration and ground truth data, broader variability in lighting and environmental factors challenges their robustness. For instance, ORB-SLAM3 (S) records an ATE of 4.80m in sequence 00, but as environmental complexity increases in sequence 02, the ATE jumps to 11.2m. This highlights SLAM’s struggle with outdoor settings where rapid changes in lighting and environment occur. On HoloSet, SLAM methods fail entirely, and they are unable to handle the unpredictable, multi-directional motion of the human-controlled XR headset. Both ORB-SLAM3 (S) and ORB-SLAM3 (M) are unable to produce valid trajectories, reflecting their difficulty in handling high frame rates and the uncontrolled dynamics of human motion. Furthermore, traditional SLAM methods are optimized for feature-rich environments; the sparse, texture-less backgrounds common in XR indoor scenes (white walls, floors) cause feature tracking to fail, leading to the inability to maintain global pose estimation. Furthermore, the higher frame rate and rolling shutter of the XR headset exacerbate feature tracking errors under fast head rotations, leading to divergence of SLAM pipelines.
@@
Deep Learning Methods. DL methods exhibit consistency across sequences but generally incur higher error rates, especially in environments with significant variability. While DL methods provide consistent performance, their ATE values are higher than hybrid methods. For instance, in MH01, DeepVO records an ATE of 1.67m – consistent but less accurate compared to KP3D ’s 0.21m. This indicates that DL methods are less precise in handling turbulent motion within controlled indoor environments. On KITTI, DL methods struggle more with outdoor complexity. For example, DeepVO cannot produce results for most sequences (denoted by “–” in the table), highlighting the difficulty of applying DL methods to rapidly changing outdoor environments, particularly when lighting conditions and scene complexity fluctuate. On the HoloSet, DL methods perform better than SLAM but still lag behind hybrid methods. For example, DeepVO records an ATE of 33.0m in suburb-jog-seq1, significantly higher than hybrid methods like KP3D. While DL methods can handle unpredictable human motion, their error rates remain higher than hybrid methods, reflecting their limitations in highly dynamic, human-centered environments. DL models lack explicit modeling of rotation and scene geometry, so fast rotations (head yaw/pitch) cause temporal inconsistencies that accumulate into trajectory errors. Additionally, HoloLens’ camera frame rate and latency differ from training data, impacting the temporal alignment of predictions.
@@ 4.5. Effect of Human Factors in XR
Deep Learning Methods. DL methods, such as DeepVO, also show an ability to track human motion better than SLAM methods but with less precision than hybrid approaches. In HoloSet’s suburb-jog-seq1, DeepVO records an ATE of 33.0m, significantly higher than KP3D ’s performance. While DL methods benefit from their inherent flexibility in handling varying motions, their higher error rates suggest they are less adept at managing rapid, unpredictable changes in orientation and environment caused by human interaction in XR scenarios. These limitations are particularly pronounced in XR motion scenarios, where abrupt head rotations introduce large angular velocities and rolling shutter distortions. DL methods typically do not incorporate inertial cues or rotation compensation, leading to temporal inconsistencies in pose estimation under such conditions. DL methods are less reliant on explicit feature tracking, giving them some robustness in texture-poor environments. However, their models are typically trained on image sequences with consistent motion patterns and fixed camera properties. In HoloSet, the combination of rolling shutter artifacts and rapid, non-repetitive head motion violates these assumptions, resulting in scale drift, temporal misalignment, and degraded accuracy compared to hybrid approaches.
@@
4.6. Quantifying the Impact of Environmental and Locomotion Challenges
Table 8. Correlation of ATE with Motion and Lighting across datasets and algorithms. Positive Motion correlation = ATE worsens with faster motion; Positive Lighting correlation = ATE worsens with brighter scenes.
Dataset Algorithm ATE vs Motion ATE vs Lighting
EuRoC ORB-SLAM3 (S) +0.24 -0.36
ORB-SLAM3 (M) -0.00 +0.19
VINS-Fusion +1.00 -0.49
DSM +0.31 -0.06
DROID-SLAM +0.28 -0.36
SfMLearner +0.21 -0.52
KP3D +0.17 +0.25
TartanVO +1.00 -0.49
DFVO +0.86 -0.86
DeepVO -0.13 +0.57
KITTI KP3D +0.16 +0.43
VINS-Fusion +1.00 +0.03
TartanVO +1.00 +0.24
HoloSet KP3D -0.03 +0.31
VINS-Fusion +1.00 +0.91
TartanVO +1.00 +0.87
Combined KP3D -0.08 +0.06
VINS-Fusion +1.00 -0.04
TartanVO +1.00 +0.02
To quantify how environmental and locomotion challenges impact tracking accuracy, we compute the Pearson correlation between per-sequence ATE and two key scene factors: Motion level (Slow=1, Medium=2, Fast=3) and Lighting level (Dark=1, Medium=2, Bright=3). We report correlations for each algorithm and dataset in Table 8. Correlations are computed using sequences where valid ATE results are available for each algorithm.
Among the evaluated methods, KP3D provides consistent results across all datasets, enabling stable correlation analysis. VINS-Fusion and TartanVO also produce valid results on multiple datasets, though in some cases with limited data. Correlation values of +1.00 in small subsets ($\leq 3$ sequences) reflect perfect alignment within the available data, but should not be interpreted as strong evidence of causal sensitivity. We report them here for completeness. Overall, KP3D exhibits a mild positive correlation with both Motion and Lighting, suggesting resilience to scene dynamics. In contrast, VINS-Fusion and TartanVO show stronger sensitivity to Motion and Lighting, particularly on human-centered HoloSet sequences, consistent with known limitations of feature-based and deep-learning-based visual odometry methods in complex environments.
Key Takeaway. This correlation analysis provides a systematic view of how scene factors, particularly motion speed and lighting conditions, impact each algorithm’s tracking performance. This deeper analysis explains why certain methods (traditional SLAM) fail under fast or low-light conditions, and why hybrid and learning-based methods remain more resilient yet still sensitive to scene dynamics. This directly addresses the need for finer-grained insights into algorithm robustness, beyond average ATE comparisons.
4.7. Computational Cost and Latency.
While our primary evaluation focuses on tracking accuracy across datasets and motion types, practical deployment in XR and IoT applications also depends critically on latency and computational cost. Table 9 summarizes the approximate computational cost and per-frame latency of each evaluated method, measured on our representative hardware platform (Intel i7-9700K CPU 3.6 GHz, 32 GB RAM, NVIDIA RTX 2070 GPU with 8 GB VRAM).
Table 9. Computational cost and latency of evaluated algorithms.
Algorithm Computational Cost Latency / FPS
ORB-SLAM3 CPU-only, heavy feature extraction + Bundle Adjustment  10–15 FPS
VINS-Fusion CPU + IMU synchronization  20 FPS
DROID-SLAM Heavy GPU (GRU + CNN backbone)  8–12 FPS
DeepVO Light GPU (CNN + RNN)  40–50 FPS
TartanVO Light GPU (CNN)  40–50 FPS
SfMLearner Light GPU (CNN)  40–50 FPS
KP3D Heavy GPU (ResNet + PoseNet + feature pipeline)  8–10 FPS
DFVO Moderate GPU (CNN + Bundle Adjustment)  20–25 FPS
DSM CPU-only, very heavy Bundle Adjustment  5–8 FPS
We observe that traditional feature-based SLAM methods (ORB-SLAM3, VINS-Fusion) remain heavily CPU-bound, with modest frame rates in the range of 10–20 FPS even without using dense representations. Deep learning-based methods (DeepVO, TartanVO, SfMLearner) achieve higher frame rates (40–50 FPS) but may lack robustness under fast motion or degraded lighting. Hybrid approaches such as KP3D and DROID-SLAM trade off accuracy for computational cost, often requiring significant GPU resources to maintain real-time operation. DSM, while accurate in specific cases, is computationally expensive Overall, our analysis highlights that latency and computational cost vary substantially across methods, and optimizing for these factors is critical for deployment on resource-constrained IoT and XR devices. We leave a more systematic evaluation of runtime efficiency under varying hardware constraints as a promising direction for future work.
4.8. Insights into Performance Limitations
Specific failure cases in SLAM and VO performance arise from the interaction between algorithmic design choices and sensing, motion, and environmental factors. For instance, traditional pipelines such as ORB-SLAM3 depend on stable feature repeatability (Campos et al., 2021) and fail in Holoset where rolling shutter distortion, low-texture indoor surfaces (Taheri and Xia, 2021), and abrupt head rotations undermine feature tracking (Pupilli and Calway, 2006). In contrast, DeepVO exhibits robustness to low-texture conditions (Jeong et al., 2021), but is highly sensitive to lighting variations (Taheri and Xia, 2021) and fast motion, since it lacks explicit geometric modeling and relies on photometric consistency. Hybrid approaches such as KP3D (Tang et al., 2020) mitigate some of these limitations by integrating learned priors with geometric constraints, but still show performance drops in high-motion or poorly lit sequences, as reflected in the correlation analysis. These patterns highlight opportunities for improving SLAM systems through adaptive strategies, which we discuss in the next section.